OpenStudy (anonymous):

Find the percent of markup. Round to the nearest tenth. $33.00 to $45.00

OpenStudy (anonymous):

to find percentage u divide the portion u need percentage of by whole thing then multiply by 100

OpenStudy (anonymous):

\[\frac{ a }{ ab} \times 100\]

OpenStudy (anonymous):

can u calculate plz

OpenStudy (anonymous):

I don't understand

OpenStudy (anonymous):

\[\frac{ 33 }{ 45 } \times 100\]

OpenStudy (anonymous):

I got a long demical

OpenStudy (anonymous):

can u plz post it

OpenStudy (anonymous):

its should be 73%